Valles Global Industries (VGI) is considering investing in a new bio-derived plastics recycling project as a means of demonstrating green solutions to their customers. Recently, VGI has been focusing its business development to address the recycling and/or reuse of plastics whenever possible. This is part of an overarching corporate goal to qualify as a “B” corporation.

Through extensive research costing approximately $500,000 over six months, they have identified three competing alternatives. Specifically, they have three options: two they can buy and one is a subcontract. Unfortunately, they have recently learned that of the two sources for processes they can buy… both are getting out of the business and only one machine of each is available. They believe their project will last five years and they will sell the machines after five years for approximately 10% of acquisition cost.

A: Wolverine Maize Machine: This machine costs $12,800,000, has annual operating costs of $85,000 and a service life of five years. Estimated site remediation is $-1,130,000 (yes, negative).

B: Hawkeye AgBusiness: This machine costs $13, 500,000, has annual operating costs of $100,000 and a service life of five years with an estimated site remediation of $-1,050,000.

C: Subcontract to Gopher Green Systems at an annual cost of $3, 500,000, guaranteed for four years and requiring a four-year commitment. This option is renewable for another four years at an increase for a maximum of $3,700,000 per year if needed because of market conditions. (That’s up another $200,000 per year). All costs are included in the subcontract.

Use a pre-tax MARR of 15% and assess options using present worth. Under what conditions might Gopher Green be a better option? Taxes and depreciation can be ignored. Discuss, in depth, your conclusions. Discuss, in detail, what happens if the project gets cancelled after two years, i.e., lifecycle ends early.
Answer #1

Wolverine Maize

CF0=-12800000

CF1= 1045000 (1130000-85000)

CF2= 1045000 (1130000-85000)

CF3=1045000 (1130000-85000)

CF4=1045000 (1130000-85000)

CF5= 2325000 (1130000-85000+1280000)

NPV(0.15,-12800000,-1150000,-1150000,-1150000,-1150000,2325000)= -7,530,966.7

Equivalent Annual Annuity = NPV * R /(1-(1+R)-N) = -7530966.7*0.15/(1-(1.15)-5) = -2,246,604.5

Hawkeye AgBusiness

CF0=-13500000

CF1= 950000 (1050000-100000)

CF2= 950000 (1050000-100000)

CF3= 950000 (1050000-100000)

CF4= 950000 (1050000-100000)

CF5= 2300000 (1050000-100000+1350000)

NPV(0.15,-13500000,950000,950000,950000,950000,2300000)= -8,386,316.6

Equivalent Annual Annuity = NPV * R /(1-(1+R)-N) = -8,386,316.6*0.15/(1-(1.15)-5) = -2,501,768.7

Gopher Green System

CF1= -3500000

CF2= -3500000

CF3= -3500000

CF4= -3500000

NPV(0.15,-3500000,-3500000,-3500000,-3500000)= -8,689,064.6

Equivalent Annual Annuity = NPV * R /(1-(1+R)-N) = -8,689,064.6*0.15/(1-(1.15)-5) = -3,043,478.3

Based on equivalent annuity Wolverine Maize Machine will be considered as it has highest EAA

IF THE PROJECT GETS CANCELLED AFTER TWO YEARS

Wolverine Maize

CF0=-12800000

CF1= 1045000 (1130000-85000)

CF2= 2325000 (1130000-85000+1280000)

NPV(0.15,-12800000,1045000,2325000)= -8,811,539.4

Hawkeye AgBusiness

CF0=-13500000

CF1= 1045000 (1050000-100000)

CF2= 2325000 (1050000-100000+1350000)

NPV(0.15,-13500000,950000,2300000)= -9,508,506.6

Gopher Green System

CF1= -3500000

CF2= -3500000

NPV(0.15,-3500000,-3500000)= -4,947,809,7

If project gets cancelled after 2 years Gopher Green project will be more feasible